OpenStudy (anonymous):

10. Input in standard form the equation of the given line. The line that passes through (-2, 4) and is parallel to x - 2y = 6

OpenStudy (ajprincess):

Solve the given equation x-2y=6 for y and can u tell me what u get?

OpenStudy (anonymous):

yes -3

OpenStudy (ajprincess):

Nope.:( what we want to do is to solve for y right? To do that we have to isolate y in thr given equation.In the given equation x has been added to x. so let us subtract x from -2y to remove it. since given one is an equation we need to do the same to the other side. so this gives x-2y-x=6-x -2y=6-x We are concerned about only y. so to get it we hav to divide it by -2 . since it has been multiplied by -2. we hav to do the same thing to the other side since this is an equation. so -2y/-2=(-x+6)/-2 y=x/2+3 compare this equation with y=mx+c and tell me the values of m and c

OpenStudy (ajprincess):

we are nt yet done.:(

OpenStudy (anonymous):

i got x-2y=-10 ty!!

OpenStudy (ajprincess):

perfect:)